Compute the rise in temperature of water ifitis brought to rest after falling over a waterfall 50.0 m high. Assume that no heat is transferred to the surroundings.


1
Expert's answer
2021-01-12T07:02:11-0500

E = m*g*h = (1 kg)*(9.81 m/s^2)*(50 m) = 490.5 J

specific heat (C) of water is 1 calorie/gram or 4.186 joule/gram

E = deltaT*C*m = 490.5 J

deltaT = (490.5 J)/(4.186 J/g)/(1000g) = 0.117 K